40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Tue May 3 18:46:06 2011, rana, Update, IOO, MC Locking not working Untitled.pngUntitled.png
    Reply  Wed May 4 11:48:01 2011, Suresh, Update, IOO, MC Locking not working 
Message ID: 4620     Entry time: Tue May 3 18:46:06 2011     Reply to this: 4621
Author: rana 
Type: Update 
Category: IOO 
Subject: MC Locking not working 

I found that the MC autolocker was OFF. Kiwamu says he turned it off because its slow. Suresh says that he has some feelings that maybe something is wrong. I'll let them describe what they know about the MC in an elog.

I checked the trend of the MC and PMC transmissions for the past 30 days:

Untitled.png

Looks like the alignment has been drifitng. PMC was corrected recently by Koji, but the alignment of the input beam to the MC or the MC itself has to be fixed. Has someone been twiddling the MC SUS alignment biases??

Attachment 1: Untitled.png  96 kB  | Hide | Hide all
Untitled.png
ELOG V3.1.3-